OpenStudy (anonymous):

I'm trying to get caught up in math . And I'm having a hard time figuring this problem out . If someone would help me it would be greatly appreciated. 5 + 3w + 2 = 5 + 2 + 3w (Points : 1) A) Commutative Property of Addition B) Commutative Property of Multiplication C) Associative Property of Addition D) Associative Property of Multiplication

OpenStudy (anonymous):

Well here is one way to go about this problem: First of all, you'll notice that there is no multiplication in this problem, however, this problem does involve addition. So you can eliminate B and D. So it's either A or C. So here is definitions from http://www.purplemath.com/modules/numbprop.htm: "The Associative Property is the rule that refers to grouping. For addition, the rule is "a + (b + c) = (a + b) + c." If you notice though, there isn't anything really grouped together... So just in case, let's look at commutative: "The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"." Since your problem involves switching around the 3w with 2, then this must be the commutative property. So your answer will be A. :) Does that makes sense?